OpenStudy (anonymous):

The defining equation of the machine f which takes a number, divides it by 5, and then adds 20 to the result is

OpenStudy (anonymous):

f(x) = x/5 + 20 f(x) = 5/x + 20 f(x) = 5 / (x+20) f(x) = x / (5+20)

OpenStudy (anonymous):

i said f(x)=x/5+20

OpenStudy (skullpatrol):

correct